OpenStudy (mony01):

Calculate the volumes of 0.20 M acetic acid, 0.20 M sodium acetate, and pure water needed to make 20.00 mL of a buffer having pH = 5.00. Determine the amounts of the acid and base that require the minimum volume of added water to reach a total volume of 20.00 mL.

OpenStudy (anonymous):

pKa of CH#COOH and CH3COONa = 4.7 pH = pKa + log salt/acid 5.00=4.7 + log CH3COONa/CH3COOH 5.00-4.7=log CH3COONa/CH3COOH 0.3 = log CH3.................. antilog 0.3 = CH3COONa / CH3COOH 1.99 =CH3COONa / CH3COOH 1.99/1 = CH3COONa / CH3COOH Its mean the ratio between salt and acid is 1.99 : 1 sum of ratio = 2.99 volume of salt from 20 mL = 1.99/2.99 X 20 = 13.3 mL volume of acid from 20 mL = 1/2.99 X 20 = 6.68 mL 13.3 mL of CH3COONa + 6.68 mL of CH3COOH = buffer of pH 5.00

OpenStudy (mony01):

would the pka still be 4.7 if instead pf pH=5, we use pH=4?